Owl Shift Summary 6/25/99

Mary_Manning (mem8f@virginia.edu)
Fri, 25 Jun 1999 07:41:02 -0500

Shift Summary for Owl Shift, 25/Jun/99

On Shift:
J. G. L.M

Hours of beam delivered at all intensities: 1.5
Hours of beam delivered at nominal intensity: 0.0
Hours of E832 data (special runs not included): 0.0
Hours of special runs (laser scans, muon runs...): 0.5

Accounting of lost time:
No beam, Resetting kumquats and starting runs

========================================================================

Runs ended this shift:
13718 - cosmics
13723 - Beam_Muon_832 4e12

Tapes brought to FCC this shift: KQ1001, ...
(including tapes for online split)
None

========================================================================

Accesses during shift:
None

========================================================================

Overview of shift: No beam for most of shift. No end-of file markers
for tapes QK*40-42 and QKB 014. Low intensity beam arrives around 5:30,
problems booting bananas and kumquats, Muon run finally started at end
of shift

========================================================================

Description of Shift:

run spill time comment
------------------------------------------------------------------------

12:00 no beam , starting TeV access, estimate 3:20
for beam
3:20 no beam, work on MI magnets, estimate 5:00
for beam
3:30 looked at track illuminations at V0 bank in B01
and B05 triggers. In B01, we see illumination
problems near counters 7 and 23 in V0. This
is not seen in B05, indicating problem
in trigger. More details in logbook

5:20 MCR calls, ready for beam

13718 5:30 Run ended

5:40 No end-of-file markers on Tapes QK* 40-42,QKB
14.
Probable cause is the drt_fresh_start prior to
run 13718
Tapes QK* 40-42 contain only cosmics, probably
can be recycled. Tape QKB 14 has laser, ped
data (probably do not want to overwrite)

5:45 Tapes QK* 43-45, QKB 15 mounted

5:50 Init aborted because an unknown process timed
out - drt_fresh_start No red buttons to indicate

error.

6:20 Process timed out again, rebooted fscc in crate
3
6:40 Process timed out again -drt_fresh_start
6:45 Reset bannanas and kumquats, finally able to
init
Found instructions in targets log.

13723 6:50 Beam_muon run, beam is low intensity, BOA ~1.9
7:00 Muon run ended, hand over to day shift

=====================================================================
Follow up Notes on Previous Shift
None

========================================================================

Items to Follow up on Next Shift

Run Beam_832_test config(doesn't have east-west up-down in trigger)